# Learning Collective Variables with Synthetic Data Augmentation through Physics-inspired Geodesic Interpolation

### v3

@@ 1 Introduction
To facilitate the efficient sampling of rare events, researchers have developed a diverse array of enhanced sampling techniques 3, which bias the simulations towards regions of interest or lower energy barriers. These techniques can roughly be divided into two classes: one that focuses on the sampling of transitions by manipulating the initial conditions4, 5, 6 and the other that introduces a bias term to the potential energy surface (PES). In this work, we concentrate on the latter class of algorithms, which encompasses a broad spectrum of methodologies, including umbrella sampling 7, metadynamics 8, and adaptive biasing force 9, 10. All of them have in common that they require the selection of one or several collective variables (CVs), special degrees of freedom that describe the rare event of interest, such that biasing along these CVs will greatly accelerate the observation of the rare transitions.
@@
Figure 1: Depiction of a common pipeline of data-driven CVs. Initially, only data from metastable states are available (left). Long and costly production runs are meaningful only when performed with a reliable CV (right). Top (previous methods): A trial CV is trained and then iteratively improved through enhanced sampling simulations, which generate more data on the transition between the metastable states. Bottom (ours): Geodesic interpolations are used to create synthetic TSE data, from which the CV can be trained in one shot, obviating the need for an iterative procedure.
@@ 2.1 Geodesic interpolation of protein conformations
Now, the interpolation between two structures $\mathbf{X},\mathbf{Y}\in\mathcal{M}$ with weights $(t,1-t)$ is given by the $w$-geodesic $\gamma_{\mathbf{X},\mathbf{Y}}^{w}:[0,1]\to\mathcal{M}$ defined as
@@
so that the interpolated structure $\gamma_{\mathbf{X},\mathbf{Y}}^{w}(t)$ is a minimizer of the inversely weighted sum of squared distances to the two structures.
Figure 2: Comparison of the unfolded–folded transition observed in a long reference trajectory and the corresponding transition generated through geodesic interpolation of the endpoints. (a) Overlay of interpolated structures (opaque) and the reference structures (transparent). The initial and final structures are identical. (b) Evolution of the donor–acceptor distances for the two key hydrogen bonds observed in the folded state, in respective colors. The dashed lines are the reference transition as a function of time, and the solid lines are the interpolated conformations as a function of parameter $t$. (c) Overlay of 5,000 interpolated samples with $t$ uniformly sampled over the range of [0, 1], displayed on the first two slow modes from TICA. The projection of the reference unbiased trajectory is shown in gray in the background. Low TIC 1 values (left side) correspond to the unfolded basin, while high values (right side) correspond to the folded basin. (d) Interpolation parameter $t$ describes the progress along the slowest mode (TIC 1), which is correlated with the transition.
@@ 2.2 Leveraging the interpolation parameter as an indicator of reaction progress
Methods based on discriminant analysis heavily rely on accurate state labels, as the loss functions are tailored to maximize the differentiation between these states. As reported in Ray et al. 24, training the CV model with reduced overlap between states can significantly enhance its performance. However, distinguishing samples metastable state samples from TSE samples becomes challenging in situations where our understanding of the system is limited. For TPI-Deep-TDA, the remedy for this problem is filtering out metastable configurations from the TSE data with the initial TDA model.24 The interpolation progress parameter $t$ can be considered as a proxy for the reaction progress, as illustrated in Figure 2. In this work, we propose training a regressor CV leveraging this additional source of information as an option to go beyond discrete state labels.
@@ 2.3 Chignolin as a model system
To demonstrate the capabilities of our enhanced sampling method for protein simulations, we utilize the chignolin variant CLN025 as a benchmark system. Despite its small size of 10 residues and 166 atoms, it folds into a stable $\beta$-hairpin structure. Simulating this folding process poses a non-trivial sampling challenge that necessitates very long simulation times in unbiased simulations. Hence, chignolin is recognized as a popular benchmark system for the evaluation of enhanced sampling methods. In this work, the 106-$\mu$s long unbiased simulation trajectory of chignolin 32 serves as reference data for assessing the accuracy and efficiency of our CVs. An average folding time of 0.6 $\mu$s was reported in the reference simulation for chignolin 32.
@@ 2.4 Transition-focused data augmentation
We begin by performing short (50 ns) MD simulations, one for each metastable state (folded and unfolded), generating 5,000 data points per state. We imposed restraints based on the Cα root-mean-squared deviation (RMSD) from the reference folded conformation, such that no transitions (folding or unfolding) could occur. Then, to mimic transitions, we perform 5,000 geodesic interpolations between the metastable state geometries using heavy atom coordinates. We control the distribution of the synthetic TSE by sampling $t$ from a target distribution. Considering the abundant data from the metastable states, we sample $t$ from a Gaussian distribution centered at 0.5 to produce more conformations closer to the transition state, as values of $t$ near 0 or 1 correspond to the samples belonging to the metastable states. As depicted in Figure 2(a), geodesic interpolation from an unfolded (left) to a folded (right) state yields structures (middle, opaque) resembling those from MD simulations (middle, transparent). Detailed information regarding the geodesic interpolations is provided in Section S1.2 of the Supporting Information (SI).
@@ 2.5 Machine learning collective variable models
Machine-learned collective variables (ML CV) are designed to map high-dimensional structural descriptors onto lower-dimensional CVs. For all models, we featurize the inputs with all 45 pairwise contacts between ten Cα atoms.
Given class labels for metastable state samples, CV learning with targeted discriminant analysis (TDA) loss, introduced by Trizio and Parrinello 20, would be a natural starting point. In the case of one-dimensional CVs, TDA assumes that there exist a non-linear transformation such that the marginal Boltzmann distribution along the CV forms a mixture of $N_{s}$ Gaussians with user defined means and variances. The TDA loss is defined as shown in eq (4), where $N_{s}$ represents the total number of states, ${\mu}_{k}$ and ${\sigma}_{k}$ are mean and standard deviation of CV model predictions for the training data with ground truth state label $k$, $\bar{\mu}_{k}$ and $\bar{\sigma}_{k}$ denote the target Gaussian mean and standard deviation corresponding to state $k$, and $\alpha$, $\beta$ are hyperparameters.
@@
We train multi-layer perceptron (MLP) models with the TDA loss, both with and without interpolated samples. The model trained solely on the two metastable states ($N_{s}=2$) is denoted as TDA, while the one trained on the two metastable states along with the synthetic transition state generated from geodesic interpolation is labeled as TDAgeo ($N_{s}=3$).
@@
More details on the model training are described in Section S1.1 in the SI.
@@ 3 Results and Discussion
Figure 4: Convergence of the free energy difference ($\Delta F$) between folded and unfolded state and the PMF at the end of the simulation as sampled with WTM-eABF using each CV. Top panel: Evolution of the $\Delta F$ estimate over the course of the trajectory. Solid lines indicate a $\pm$ 4 kJ/mol range (chemical accuracy) around the $\Delta F$ value obtained from the long unbiased reference runs (dashed line). Bottom panel: Comparison of reference PMF obtained by projection of reference data (dotted line) and those obtained from 1 $\mu$s WTM-eABF simulations with the given CV. The shaded areas represent the standard deviation of the independent simulations, and the solid line is the mean.
@@ 3.1 Geodesic interpolation resembles state-state transitions
In this section, we investigate the properties of the interpolated data and the parameter $t$. The intermediate structures generated using geodesic interpolation smoothly interpolate between the unfolded and folded structures, with a close resemblance to the unbiased reference trajectory over simulation time, as visualized in Fig 2(a). The donor–acceptor distances of the two key hydrogen bonds between Asp3–Thr6 and Asp3–Thr8 are also smoothly interpolated, following the trend of those from the unbiased trajectory (Figure 2(b)).
@@
Time-lagged independent component analysis (TICA) is a dimensionality reduction technique that effectively diagonalizes a time-lagged correlation matrix 40, 41. Where principal component analysis (PCA) returns eigenvectors that are linear combinations of the input features that maximize the corresponding variance, TICA returns the combinations with the slowest motions. The first time-lagged independent component (TIC) represents the slowest collective motions of the chignolin system, capturing the unfolded–folded transition, and the second component resolves the three clusters within the folded state. Figure 2(c) shows that the interpolated structures and the reference transition paths have very similar projections on the first two TICs. Furthermore, as shown in Figure 2(d), the interpolation parameter $t$ has a strong correlation with the slowest mode of the system (TIC 1), indicating that $t$ is in fact a useful descriptor of the progress of chignolin folding. Finally, the comparison of the $t$ and $\hat{t}$ in Figure 3(c) demonstrates that the two values are well aligned with each other. This indicates that the endpoint dependence of the parameter $t$ is weak, allowing it to be interpreted as a transition progress between two metastable state ensembles rather than being linked to specific samples of them.
@@
For chignolin, interpolating 5,000 frames took approximately 40 minutes, which is about half the time of each 50 ns unbiased metastable state simulation needed to extract the same number of frames. Further details regarding computation time and scaling results are provided in the SI, Section S3.
@@
Compared to the physical TSE, such as those derived from the on-the-fly probability enhanced sampling (OPES) flooding approach 42, 24, the pseudo-TSE generated from the geodesic interpolation is advantageous in being simulation parameter-free, faster, and allowing for controllable generation (by means of $t$) and facilitating generation of decorrelated samples. However, it is purely geometric in nature and is not guaranteed to represent the physical TSE. On the other hand, the generation approach utilizing enhanced sampling simulations produces physically realistic structures, but is slower and requires the system to escape the metastable basin. It also results in correlated samples, i.e., one run only produces a single path exploring the TSE, which then quickly goes over into the other metastable basin. Additionally, it necessitates some simulation parameter settings, such as the expected barrier height. Despite these differences, both methods share the benefit of being parallelizable, enabling efficient computation across multiple processors. We expect that the choice of method may depend on the user’s familiarity with the enhanced sampling protocol and the available computational budget for simulations.
@@ 3.2 Data augmentation improves ML CV
For each of the statistically independent 1-$\mu$s WTM-eABF enhanced sampling simulations, with the first 100 ns discarded as equilibration time, we obtained the PMF and calculated the free energy of folding ($\Delta F$, eq (5)) as a function of simulation time. The convergence of $\Delta F$ and the final PMFs obtained from the enhanced sampling runs are reported in Figure 4. We further compare the sampled PMFs with those obtained from the reference trajectory by investigating the difference between $\Delta F$ and the reference value $\Delta F_{\text{ref}}$ and evaluating the PMF MAE (eq (6)), collected in Table 1.
@@
Figure 4 and Table 1 show that models trained with data from the meta-stable states augmented with filtered pseudo-TSE data perform the best (TDAgeo (filtered) and Reggeo (filtered)). They are the only models that, on average, successfully approach the reference value of $\Delta F$ within the chemical accuracy threshold of 4 kJ/mol. Surprisingly, all other models predict the unfolded state to be more stable. Further, using all 5,000 pseudo-TSE configurations to train a three-state TDA model, TDAgeo (unfiltered), shows no improvement compared to the two-state TDA model. The reason for this likely being the significant overlap of the distribution of structures generated by the geodesic interpolation and those from the meta-stable state simulations. Both TDAgeo and Reggeo show faster convergence in terms of the average $\Delta F$ value, even if it is to the wrong value, than TDA without data augmentation. Notably, the CVs trained on the filtered data quickly converge to the reference $\Delta F$ in less than 500 ns and demonstrate a very good reproduction of the reference PMF.
@@ 3.3 Analysis of CV model predictions for sampled configurations
To better understand the behavior of the different ML CVs, we show two low-dimensional projections of the data. Firstly, the TICA analysis was performed on the long, unbiased reference simulation. In Figure 5, we have projected all conformations of the reference trajectory, as well as the trajectories generated from the different enhanced sampling runs, onto the space spanned by the first two TICs (TIC 1 and TIC 2). The points were colored by scaled CV values (such that each CV is zero for the unfolded and one for the folded minimum on the PMF). Secondly, in Figure 6 we plotted the CV value of each configuration against its committor estimate, where the committor function was obtained in a data-driven manner from the recent study by Kang et al. 27. The points are colored by the number of hydrogen bonds formed in the native folded conformation, a physical observable that correlates well with the folding process of chignolin 43, 44. A detailed illustration of the hydrogen bond formation in the chignolin folding process can be found in Section S1.3 of the SI.
@@
The first row of Figure 6 shows the TICA projection of the unbiased reference data, while the bottom row shows the data obtained from the respective enhanced sampling runs, similar to Figure 5. Figure 6 shows that the number of formed hydrogen bonds (as indicated by the color of the dots) correlates well with the committor value; however, for all models and all runs, there are configurations with no formed contacts but committor values indicating an almost folded state. Note that we have used one minus the committor value from Kang et al. 27 such that it is zero for unfolded and one for folded conformations. The upper row shows the expected sigmoidal curves for committor vs. CV, as the ML CVs change more gradually. All CV models show a clear separation of the extreme committor values (near zero and one) for the unbiased geometries, which is in agreement with the first row of Figure 5 and explains the general agreement for the reference $\Delta F$ among all ML CV models. However, one can observe a distinct broadening, especially for all TDA-type models in the bottom row. In order to observe a good estimation of the reaction free energy, a CV needs to clearly distinguish between the different meta-stable state, i.e., it needs to separate committor values of 0 and 1. Especially for the unfiltered TDAgeo and Reggeo this is not the case. For Reggeo (unfiltered), one can see the overlap of extreme committor values for predicted CV values of around 0.5, which explains the noticeable lowering of the PMF in the transition region for this model in Figure 4. The bottom row of Figure 6 clearly explains why the two models trained on the TDA-filtered data perform better than those trained on unfiltered data.

### v4

@@ 1 Introduction
To facilitate the efficient sampling of rare events, researchers have developed a diverse array of enhanced sampling techniques 3, which bias the simulations toward regions of interest or lower energy barriers. These techniques can roughly be divided into two classes: one that focuses on the sampling of transitions by manipulating the initial conditions4, 5, 6 and the other that introduces a bias term to the potential energy surface (PES). In this work, we concentrate on the latter class of algorithms, which encompasses a broad spectrum of methodologies, including umbrella sampling 7, metadynamics 8, and adaptive biasing force 9, 10. All of them have in common that they require the selection of one or several collective variables (CVs), special degrees of freedom that describe the rare event of interest, such that biasing along these CVs will greatly accelerate the observation of the rare transitions.
@@
Figure 1: Depiction of a common pipeline for data-driven CVs. Initially, only data from metastable states are available (left). Long and costly production runs are meaningful only when performed with a reliable CV (right). Top (previous methods): A trial CV is trained and then iteratively improved through enhanced sampling simulations, which generate more data of the transition between the metastable states. Bottom (ours): Geodesic interpolations are used to create synthetic TSE data, from which the CV can be trained in one shot, obviating the need for an iterative procedure.
@@ 2.1 Geodesic interpolation of protein conformations
Now, the interpolation $\mathbf{Z}$ between two structures $\mathbf{X},\mathbf{Y}\in\mathcal{M}$ with weights $(t,1-t)$ is given by the $w$-geodesic $\gamma_{\mathbf{X},\mathbf{Y}}^{w}:[0,1]\to\mathcal{M}$ defined as
@@
so that the interpolated structure $\gamma_{\mathbf{X},\mathbf{Y}}^{w}(t)$ is a minimizer of the inversely weighted sum of squared distances to the two structures. See Figure 2(a) for an illustrative example of interpolated structures.
Figure 2: Comparison of the unfolded–folded transition observed in a long reference trajectory and the corresponding transition generated through geodesic interpolation of the end points. (a) Overlay of interpolated structures (opaque) and the reference structures (transparent). The initial and final structures are identical. (b) Evolution of the donor–acceptor distances for the two key hydrogen bonds observed in the folded state, in respective colors. The dashed lines are the reference transition as a function of time, and the solid lines are the interpolated conformations as a function of parameter $t$. (c) Overlay of 5,000 interpolated samples with $t$ uniformly sampled over the range of [0, 1], displayed on the first two slow modes from TICA. The projection of the reference unbiased trajectory is shown in gray in the background. Low TIC 1 values (left side) correspond to the unfolded basin, while high values (right side) correspond to the folded basin. (d) Interpolation parameter $t$ correlates with the progress along the slowest mode (TIC 1), which describes the transition.
@@ 2.2 Leveraging the interpolation parameter as an indicator of reaction progress
Methods based on discriminant analysis heavily rely on accurate state labels, as the loss functions are tailored to maximize the differentiation between these states. As reported in Ray et al. 24, training the CV model with reduced overlap between states can significantly enhance its performance. However, distinguishing metastable state samples from TSE samples becomes challenging in situations where our understanding of the system is limited. For TPI-Deep-TDA, the remedy for this problem is filtering out metastable configurations from the TSE data with the initial TDA model.24 The interpolation progress parameter $t$ can be considered as a proxy for the reaction progress, as illustrated in Figure 2. In this work, we propose training a regressor CV leveraging this additional source of information as an option to go beyond discrete state labels.
@@ 2.3 Chignolin as a model system
To demonstrate the capabilities of our enhanced sampling method for protein simulations, we utilize the chignolin variant CLN025 as a benchmark system. Despite its small size of 10 residues and 166 atoms, it folds into a stable $\beta$-hairpin structure. Simulating this folding process poses a nontrivial sampling challenge that necessitates very long simulation times in unbiased simulations. Hence, chignolin is recognized as a popular benchmark system for the evaluation of enhanced sampling methods. In this work, the 106-$\mu$s long unbiased simulation trajectory of chignolin 32 serves as reference data for assessing the accuracy and efficiency of our CVs. An average folding time of 0.6 $\mu$s was reported in the reference simulation for chignolin 32.
@@ 2.4 Transition-focused data augmentation
We begin by performing short (50 ns) MD simulations, one for each metastable state (folded and unfolded), generating 5,000 data points per state. We imposed restraints based on the Cα root-mean-squared deviation (RMSD) from the reference folded conformation, such that no transitions (folding or unfolding) can occur. Note that, while the TPI-Deep-TDA study 24 sampled configurations every 1 ps, we sampled every 10 ps, corresponding to a 10-fold reduction in the number of metastable state data points compared to their study. In our opinion, closer temporal spacing between the sampled geometries would not improve the performance of machine learning collective variables (ML CVs) due to the similarity of the geometries. Then, to mimic transitions, we perform 5,000 geodesic interpolations between the metastable state geometries using heavy atom coordinates. We control the distribution of the synthetic TSE by sampling $t$ from a target distribution. Considering the abundant data from the metastable states, we sample $t$ from a Gaussian distribution centered at 0.5 to produce more conformations closer to the transition state, as values of $t$ near 0 or 1 correspond to the samples belonging to the metastable states. As depicted in Figure 2(a), geodesic interpolation from an unfolded (left) to a folded (right) state yields structures (middle, opaque) resembling those from MD simulations (middle, transparent). Detailed information regarding the geodesic interpolations is provided in Section S1.2 of the Supporting Information (SI).
@@ 2.5 Machine learning collective variable models
ML CVs are designed to map high-dimensional structural descriptors onto lower-dimensional CVs. For all models, we featurize the inputs with all 45 pairwise contacts between ten Cα atoms.
Given class labels for metastable state samples, CV learning with targeted discriminant analysis (TDA) loss, introduced by Trizio and Parrinello 20, would be a natural starting point. In the case of one-dimensional CVs, TDA assumes that there exist a nonlinear transformation such that the marginal Boltzmann distribution along the CV forms a mixture of $N_{s}$ Gaussians with user defined means and variances. The TDA loss is defined as shown in eq (4), where $N_{s}$ represents the total number of states, ${\mu}_{k}$ and ${\sigma}_{k}$ are mean and standard deviation of CV model predictions for the training data with ground truth state label $k$, $\bar{\mu}_{k}$ and $\bar{\sigma}_{k}$ denote the target Gaussian mean and standard deviation corresponding to state $k$, and $\alpha$, $\beta$ are hyperparameters.
@@
We train multilayer perceptron (MLP) models with the TDA loss, both with and without interpolated samples. The model trained solely on the two metastable states ($N_{s}=2$) is denoted as TDA, while the one trained on the two metastable states along with the synthetic transition state generated from geodesic interpolation is labeled as TDAgeo ($N_{s}=3$).
@@
More details on the model training are described in SI, Section S1.1.
@@ 3 Results and Discussion
Figure 4: Convergence of the free energy difference ($\Delta F$) between folded and unfolded state and the PMF at the end of the simulation as sampled with WTM-eABF using each CV. Top panel: Evolution of the $\Delta F$ estimate over the course of the trajectory. Horizontal solid lines indicate a $\pm$ 4 kJ/mol range (chemical accuracy) around the $\Delta F$ value obtained from the long unbiased reference runs (dashed line). Bottom panel: Comparison of reference PMF obtained by projection of reference data (dotted line) and those obtained from 1 $\mu$s WTM-eABF simulations with the given CV. The shaded areas represent the standard deviation of the independent simulations, and the solid line is the mean.
@@ 3.1 Geodesic interpolation resembles state-state transitions
In this section, we investigate the properties of the interpolated data and the parameter $t$. The intermediate structures generated by using geodesic interpolation smoothly interpolate between the unfolded and folded structures, with a close resemblance to the unbiased reference trajectory over simulation time, as visualized in Figure 2(a). The donor–acceptor distances of the two key hydrogen bonds between Asp3–Thr6 and Asp3–Thr8 are also smoothly interpolated, following the trend of those from the unbiased trajectory (Figure 2(b)).
@@
Time-lagged independent component analysis (TICA) is a dimensionality reduction technique that effectively diagonalizes a time-lagged correlation matrix 40, 41. Where principal component analysis (PCA) returns eigenvectors that are linear combinations of the input features that maximize the corresponding variance, TICA returns the combinations with the slowest motions. The first time-lagged independent component (TIC) represents the slowest collective motions of the chignolin system, capturing the unfolded–folded transition, and the second component resolves the three clusters within the folded state. Figure 2(c) shows that the interpolated structures and the reference transition paths have very similar projections on the first two TICs. Furthermore, as shown in Figure 2(d), the interpolation parameter $t$ has a strong correlation with the slowest mode of the system (TIC 1), indicating that $t$ is in fact a useful descriptor of the progress of chignolin folding. Finally, the comparison of $t$ and $\hat{t}$ in Figure 3(c) demonstrates that the two values are well aligned with each other. This indicates that the end point dependence of the parameter $t$ is weak, allowing it to be interpreted as a measure of the transition progress between two metastable state ensembles rather than being linked to specific samples of them.
@@
For chignolin, interpolating 5,000 frames took approximately 40 min, which is about half the time of each 50 ns unbiased metastable state simulation needed to extract the same number of frames. Further details regarding computation time and scaling results are provided in the SI, Section S3.
@@
Compared to the physical TSE, such as those derived from the on-the-fly probability enhanced sampling (OPES) flooding approach 42, 24, the pseudo-TSE generated from the geodesic interpolation is advantageous in being simulation parameter-free, faster, and allowing for controllable generation (by means of $t$) and facilitating the generation of decorrelated samples. However, this pseudo-TSE is purely geometric in nature and is not guaranteed to represent the physical one. On the other hand, utilizing enhanced sampling simulations to generate samples produces physically realistic structures, but is slower and requires the system to escape the metastable basin. It also results in correlated samples, i.e., one run only produces a single path exploring the TSE, which then quickly goes over into the other metastable basin. Additionally, it necessitates some simulation parameter settings, such as the expected barrier height. Despite these differences, both methods share the benefit of being parallelizable, enabling efficient computation across multiple processors. We expect that the choice of method may depend on the user’s familiarity with the enhanced sampling protocol and the available computational budget for simulations.
@@ 3.2 Data augmentation improves ML CV
For each of the statistically independent 1-$\mu$s WTM-eABF enhanced sampling simulations, with the first 100 ns discarded as equilibration time, we obtained the PMF and calculated the free energy of folding ($\Delta F$, eq (5)) as a function of simulation time. The convergence of $\Delta F$ and the final PMFs obtained from the enhanced sampling runs are reported in Figure 4. We further compare the sampled PMFs with those obtained from the reference trajectory by investigating the difference between $\Delta F$ and the reference value $\Delta F_{\text{ref}}$ and evaluating the PMF MAE (eq (6)), as collected in Table 1.
@@
Figure 4 and Table 1 show that models trained with data from the metastable states augmented with filtered pseudo-TSE data perform the best (TDAgeo (filtered) and Reggeo (filtered)). They are the only models that, on average, successfully approach the reference value of $\Delta F$ within the chemical accuracy threshold of 4 kJ/mol. Surprisingly, all other models predict the unfolded state to be more stable. Further, using all 5,000 pseudo-TSE configurations to train a three-state TDA model, TDAgeo (unfiltered), shows no improvement compared to the two-state TDA model. The reason for this likely being the significant overlap of the distribution of structures generated by the geodesic interpolation and those from the metastable state simulations. Both TDAgeo and Reggeo show faster convergence in terms of the average $\Delta F$ value, even if it is to the wrong value, than TDA without data augmentation. Notably, the CVs trained on the filtered data quickly converge to the reference $\Delta F$ in less than 500 ns and demonstrate a very good reproduction of the reference PMF.
@@ 3.3 Analysis of CV model predictions for sampled configurations
To better understand the behavior of the different ML CVs, we show two low-dimensional projections of the data: TIC 1–TIC 2 (Figure 5) and CV value–committor prediction (Figure 6). First, in Figure 5, we have projected all conformations of the reference trajectory, as well as the trajectories generated by the different enhanced sampling runs, onto the space spanned by the first two TICs (TIC 1 and TIC 2). The TICA analysis was only performed on the long, unbiased reference simulation, i.e., the projection operation is the same for all subplots, and only the data varies. The points were colored by scaled CV values (such that each CV is zero for the unfolded and one for the folded minimum on the PMF). Second, in Figure 6, we plotted the CV value of each configuration against its committor estimate, where the committor function was obtained in a data-driven manner in the recent study by Kang et al. 27. The points are colored by the number of hydrogen bonds formed in the native folded conformation, a physical observable that correlates well with the folding process of chignolin 43, 44. A detailed illustration of the hydrogen bond formation in the chignolin folding process can be found in SI, Section S1.3.
@@
The first row of Figure 6 shows the TICA projection of the unbiased reference data, while the bottom row shows the data obtained from the respective enhanced sampling runs, similar to Figure 5. Figure 6 shows that the number of formed hydrogen bonds (as indicated by the color of the dots) correlates well with the committor value; however, for all models and all runs, there are configurations with no formed contacts but committor values indicating an almost folded state. Note that we have used one minus the committor value from Kang et al. 27 such that it is zero for unfolded and one for folded conformations. The upper row shows the expected sigmoidal curves for committor vs. CV, as the ML CVs change more gradually. For the unbiased geometries, all CV models show a clear separation of the extreme committor values (near zero and one), which is in agreement with the first row of Figure 5 and explains the general agreement for the reference $\Delta F$ among all ML CV models. However, one can observe a distinct broadening, especially for all TDA-type models in the bottom row of Figure 6. In order to observe a good estimation of the reaction free energy, a CV needs to clearly distinguish between the different metastable state, i.e., it needs to separate committor values of 0 and 1. Especially for the unfiltered TDAgeo and Reggeo this is not the case. For Reggeo (unfiltered), one can see the overlap of extreme committor values for predicted CV values of around 0.5, which explains the noticeable lowering of the PMF in the transition region for this model in Figure 4. The bottom row of Figure 6 clearly explains why the two models trained on the TDA-filtered data perform better than those trained on unfiltered data.